# An Optimization Framework to Enforce Multi-View Consistency for Texturing 3D Meshes Using Pre-Trained Text-to-Image Models

### v1

@@ Abstract
A fundamental problem in the texturing of 3D meshes using pre-trained text-to-image models is to ensure multi-view consistency. State-of-the-art approaches typically use diffusion models to aggregate multi-view inputs, where common issues are the blurriness caused by the averaging operation in the aggregation step or inconsistencies in local features. This paper introduces an optimization framework that proceeds in four stages to achieve multi-view consistency. Specifically, the first stage generates an over-complete set of 2D textures from a predefined set of viewpoints using an MV-consistent diffusion process. The second stage selects a subset of views that are mutually consistent while covering the underlying 3D model. We show how to achieve this goal by solving semi-definite programs. The third stage performs non-rigid alignment to align the selected views across overlapping regions. The fourth stage solves an MRF problem to associate each mesh face with a selected view. In particular, the third and fourth stages are iterated, with the cuts obtained in the fourth stage encouraging non-rigid alignment in the third stage to focus on regions close to the cuts. Experimental results show that our approach significantly outperforms baseline approaches both qualitatively and quantitatively.
@@ 1 Introduction
Specifically, our approach proceeds in four stages. The first stage uses a pre-trained text-to-image model to generate an overcomplete set of candidate textures from a predefined set of 2D viewpoints. The second stage then optimizes a subset of views that are mutually consistent. The third stage performs joint alignment among selected 2D textures to maximize consistency across overlapping regions. The fourth stage cuts and stitches the aligned texture images into a consistently textured mesh output. The third and fourth stages are iterated in an alternating manner, where the cuts obtained in the fourth stage guide the non-rigid alignment procedure so that alignment focuses on regions near the optimized cuts.
Experimental results show that our approach outperforms state-of-the-art text-to-texture approaches both qualitatively and quantitatively. Qualitatively, our approach is much more photorealistic and does not present discontinuous or blurry texture elements that are frequent among other approaches. Quantitatively, our approach outperforms the state-of-the-art approach in the FID score.
@@ 2 Related Work
A challenge in text-to-textured-mesh models is ensuring the multi-view consistency among the text-to-image outputs. Dreamfusion [31] optimizes the geometry and color of a NeRF through SDS loss. Zero123 [24] uses an image and a camera transformation as a condition to synthesize novel views, which cannot handle large differences in pixels caused by extreme relative transformations. TextMesh [47] instead uses four canonical viewpoints to render a mesh with an overall satisfactory geometric appearance despite the blurriness. Several methods  [50, 52, 40, 25, 51, 39] are trained on 3D rendered datasets [7] to enhance consistency but tend to lose realistic details in generated images. Moreover, most of them are agnostic to depth information, which is crucial for controlling image quality.
Text-to-Texture Models.
@@
Several pioneers  [29, 30, 4] are trained on a semantic CLIP loss, but struggle to get image and text semantic alignment and generate high quality textures. Trained with a large dataset [36], diffusion models [35] can generate high-quality images with rich details from text prompts. Additional conditions, such as depth maps and inpainting masks [56], can provide 3D structure-aware on 2D through rendering. Sequential methods [2, 34] generate 2D images on these conditions and incrementally merge them into a partially textured mesh, However, this approach can easily lead to suboptimal results due to accumulated errors and a lack of global perception. Using a latent UV map [28, 1, 18, 11, 26] as a unified representation has been shown to increase global consistency. Since latent pixels compress larger region of RGB images, the decoded RGB pixels cannot guarantee 3D consistency, which leads to the blurriness or inconsistent details in the textures. Some methods [54, 55] finetune a diffusion model on latent uv space to avoid projection, but also limited to the quality of uv map. Other methods [8, 53, 3] focus on PRB material generation. In contrast, our approach performs alternating optimization of non-rigid alignment among the selected images, as well as joint optimization of the best cuts to guide the non-rigid alignment.
View Selection.
@@
The quality of a multi-view representation is highly dependent on the distribution of source 2D views. View selection is related to the view saliency problem, which has been extensively studied in the existing literature [21, 9, 5, 37, 22, 17, 42]. Closely related to our work, Sun et al. [44] present an approach to optimize camera poses to maximize the coverage of an input scene. Wei et al. [49] optimize viewpoints to maximize the quality of fusion of the learned 2D features. We introduce a view selection approach that selects a subset of consistent views. Driven by the goal of enforcing multi-view consistency, we propose a novel formulation that optimizes a view-consistent score among the chosen images, while constraining that the selected images provide complete coverage of the underlying 3D model.
Texture Cutting and Stitching.
@@
Cutting and stitching overlapping images is a fundamental approach to texture synthesis. The pioneering work of Efros and Freeman [10] optimizes cuts between pairs of overlapping patches for texture synthesis. Our method formulates cut optimization and texture stitching as a second-order MRF problem on the mesh faces, where the second-order term models view consistency across different images. This MRF formulation is introduced by Waechter et al. [48], and our contribution lies in the jointly aligned images as input. In particular, we find it essential to alternate between solving the MRF problem and using the MRF output to refine its inputs.
@@ 3.2 Approach Overview
View selection. The second stage selects a subset of RGB-D images that are consistent with each other. This is formulated as a solution to a quadratic assignment problem in which the objective function models the consistency scores between pairs of RGB-D images, under the constraint that the selected images provide a complete coverage of $\mathcal{M}$. We introduce a novel sequential semidefinite programming approach to solve this view selection problem accurately and efficiently.
View alignment. The third stage performs joint non-rigid warpping of the selected images so that the pixel colors after warpping agree with each other. We first solve a joint optimization problem to alleviate illumination changes in the selected images. We then use SIFTFlow [23] to precompute dense correspondences between overlapping views. The resulting correspondences are fed into another joint optimization problem to optimize the deformation of each view so that the deformed images are view-consistent.
@@ 4 Approach
As illustrated in Figure 2, our approach consists of four stages. We first describe view generation and view selection in Section 4.1 and Section 4.2, respectively. We then elaborate on view alignment in Section 4.3. Finally, we introduce texture cutting and stitching and discuss how to alternate between this stage and view alignment in Section 4.4.
@@ 4.1 Stage I: View Generation
The second phase, which runs between diffusion steps 800-200, adopts latent-uv [28], a popular data-free approach to enforce multi-view consistency. The key idea is to aggregate multiple views in a shared latent map that is then mapped back to individual views. Although improved consistency is achieved, image details can still be dropped during the aggregation step, which performs some type of averaging, as shown in Figure 4. Therefore, similar to SyncMVD [26], we do not perform latent-uv to generate the final output.
Figure 4: The left image is the original RGB image. The second column is the warping image caused by latent UV. The third column is the warping image by RGB UV. The straight lines of the warping image by latent UV become a set of horizontal line segments, and the details are blurred.
@@
As shown in Figure 3, the multi-view image generation results are consistent in global style and shape. However, local shapes and styles may differ and there are also drifts. Another important idea is to generate an over-complete set of images and operate an a subset of more consistent images to generate the output. This approach nicely mitigates inconsistency introduced in the third phase, which is performed on each view independently.
@@ 4.2 Stage II: View Selection
$\underset{\boldsymbol{z}}{\textup{minimize}}\quad\boldsymbol{s}^{T}\boldsymbol{z}+\lambda\boldsymbol{z}^{T}S\boldsymbol{z}\qquad\textup{subject to}\quad C\boldsymbol{z}\geq\boldsymbol{1}.$ (1)
Equation (1) is hard to optimize, due to non-convexity and integer variables. We introduce two relaxations. The first relaxes $z_{i}$ so that they can take real values between 00 and $1$, and the second uses the semidefinite programming relaxation to $\boldsymbol{z}^{S}\boldsymbol{z}$. More precisely, we introduce an additional matrix variable $Y=\boldsymbol{z}\boldsymbol{z}^{T}$ and apply the standard relaxation approach as $Y\succeq\boldsymbol{z}\boldsymbol{z}^{T}$ and $\textup{diag}(Y)=\boldsymbol{z}$. Furthermore, we maintain an index set $\mathcal{I}_{f}$ that specifies all currently selected images. Initially, we $\mathcal{I}_{f}=\emptyset$. This leads to the following semidefinite program:
$\underset{\boldsymbol{z},Y}{\textup{minimize}}$ $\quad\langle\boldsymbol{s},\boldsymbol{z}\rangle+\lambda\langle S,Y\rangle$
@@
where $Y\succeq\boldsymbol{z}\boldsymbol{z}^{T}$ is identical to (4). We then augment $\mathcal{I}_{f}$ with $i^{\star}=\textup{argmax}_{i\notin\mathcal{I}_{f}}z_{i}$. This procedure is iterated until $C\boldsymbol{z}\geq 1$ is satisfied. For simplicity, we assume that the selected views are the first $k<n$ images $I_{i},1\leq i\leq k$.
Image score $s(I)$. Our key observation is that if the quality of an image $I$ is high, then the pixel colors tend to repeat among other input images. This motivates us to build a probabilistic color model for each vertex $v\in\mathcal{V}$ (cf., Figure 5 (a)). Specifically, denote $\mathcal{I}_{v}\subset\mathcal{I}$ as the subset of images in which $v$ is visible. Let $\boldsymbol{c}(v,I)$ be the color of $v$ in $I$. We first perform mean shift clustering [6] to extract the largest cluster of $\{\boldsymbol{c}(v,I)|I\in\mathcal{I}_{v}\}$, which represents the most represented color. With $\boldsymbol{c}_{v}$ and $\sigma_{v}$ we denote the center and variance of this cluster. We then define
@@
Given an image $I$, let $\mathcal{V}_{I}\subset\mathcal{V}$ collect all visible vertices in $I$. We sort the values $\{p(v,\boldsymbol{c}(v,I))|v\in\mathcal{V}_{I}\}$ in increasing order and set the score $s^{\prime}(I)$ of $I$ as the $t\%$ percentile ($t=20$) of $p(v,\boldsymbol{c}(v,I))$. The final score is given by $s(I)=(s_{\max}-s^{\prime}(I))/(s_{\max}-s_{\min})$ where $s_{\max}(s_{\min})=\max_{I}(\min_{I})s^{\prime}(I)$. As shown in Figure 5 (Left), high-quality images tend to have small values in $s(I)$.
Consistency score $s(I_{i},I_{j})$. For the purpose of view selection, we find that defining $s(I_{i},I_{j})$ based on mean color differences without alignments is sufficient. We set $s(I_{i},I_{j})=0$ if $I_{i}$ and $I_{j}$ do not overlap. Figure 5 (Right) shows that the consistency score is critical to selecting consistent images.
@@ 4.3 Stage III: View Alignment
The third stage of our approach performs non-rigid alignment among the selected RGB-D images, making them consistent among their overlapping regions. As illustrated by Figure 6, we achieve this goal in two steps. The first step adjusts the pixel colors together to reduce the illumination variations in the selected images. The second step performs joint image warpping to align image features across the selected images. In both steps, we parameterize the deformation of each image using the free-form deformation (FFD) [38] on a regular $L\times L$ lattice (we choose $L=20$ for color adjustment and $L=40$ for image warpping). Note that it is possible to avoid illumination variations by training text-to-image models from rendered 3D shapes. However, our goal is to use photorealistic images from pre-trained text-to-image models. Therefore, we chose to solve an optimization problem to consistently reduce color variations.
@@
$\hat{I}_{i}^{x}(\boldsymbol{p})=B(\boldsymbol{p})\boldsymbol{c}_{i,x}^{1}+I_{i}^{x}B(\boldsymbol{p})\big{(}\boldsymbol{1}-\boldsymbol{c}_{i,x}^{0}-\boldsymbol{c}_{i,x}^{1}\big{)},$ (6)
@@
$\min_{\{\boldsymbol{c}_{i,x}^{l}\}}$ $\sum\limits_{x\in\{r,g,b\}}\Big{(}\sum\limits_{i\neq j}\sum\limits_{\boldsymbol{p}\in I_{ij}}\|I_{i}^{x}(\boldsymbol{p})-I_{j}^{x}(g_{ij}(\boldsymbol{p}))\|^{2}+\mu\sum\limits_{i=1}^{k}\sum\limits_{l=0}^{1}{\boldsymbol{c}_{i,x}^{l}}^{T}L{\boldsymbol{c}_{i,x}^{l}}\Big{)}$
@@
Figure 6: Illustration of the view alignment stage. We show the result of two overlapping images through the joint view alignment procedure. (a) Two input images. (b) After adjustment of the global illumination. (c) From left to right: overlapping regions of each input image, warpped images using ground-truth maps between the overlapping regions, overlaid results of the first two columns, and overlaid after SIFTFlow alignment. (d) Overlaid alignments after joint alignment. The pairwise SIFTFlow alignments are preserved.
Image warpping. Let $(O_{ij}\subset I_{ij},O_{ji}\subset I_{ji})$ be the focused overlapping subsets between $I_{i}$ and $I_{j}$. Our goal is to deform each $I_{i}$ so that $O_{ij}$ and $O_{ji}$ are aligned among the deformed images. Initially, $O_{ij}=I_{ij}$ and $O_{ji}=I_{ji}$. After determining the optimal cut between $I_{i}$ and $I_{j}$, we recompute $O_{ij}$ and $O_{ji}$ as the dilated regions along the cuts. This allows us to focus on alignments among potentially more consistent regions to improve alignment results.
Let $f_{ij}:g_{ij}(O_{ji})\rightarrow O_{ij}$ and $f_{ji}:g_{ji}(O_{ij})\rightarrow O_{ji}$ be dense pixel-wise alignments between the focused overlapping regions under the view of each image. In our experiments, we compute $f_{ij}$ and $f_{ji}$ using SIFTFlow [23], a state-of-the-art method that establishes high-fidelity dense correspondences (Figure 6 (c)).
@@
where $\mu_{1}=2$ and $\mu_{2}=200$ in all our experiments. Since (8) consists of non-linear least squares, we apply the Gauss-Newton method for optimization.

### v2

@@ Abstract
A fundamental problem in the texturing of 3D meshes using pre-trained text-to-image models is to ensure multi-view consistency. State-of-the-art approaches typically use diffusion models to aggregate multi-view inputs, where common issues are the blurriness caused by the averaging operation in the aggregation step or inconsistencies in local features. This paper introduces an optimization framework that proceeds in four stages to achieve multi-view consistency. Specifically, the first stage generates an over-complete set of 2D textures from a predefined set of viewpoints using an MV-consistent diffusion process. The second stage selects a subset of views that are mutually consistent while covering the underlying 3D model. We show how to achieve this goal by solving semi-definite programs. The third stage performs non-rigid alignment to align the selected views across overlapping regions. The fourth stage solves an MRF problem to associate each mesh face with a selected view. In particular, the third and fourth stages are iterated, with the cuts obtained in the fourth stage encouraging non-rigid alignment in the third stage to focus on regions close to the cuts. Experimental results show that our approach significantly outperforms baseline approaches both qualitatively and quantitatively. Project page: https://aigc3d.github.io/ConsistenTex.
Keywords:
3D texturing Multi-view consistency Optimization
22footnotetext: Corresponding authors: list.dzl@alibaba-inc.com, huangqx@utexas.edu.
@@ 1 Introduction
Specifically, our approach proceeds in four stages. The first stage uses a pre-trained text-to-image model to generate an overcomplete set of candidate textures from a predefined set of viewpoints. The second stage optimizes a subset of views that are mutually consistent. The third stage performs joint alignment among selected 2D textures to maximize consistency across overlapping regions. The fourth stage cuts and stitches the aligned texture images into a consistently textured mesh. The third and fourth stages are iterated in an alternating manner, where the cuts obtained in the fourth stage guide the non-rigid alignment procedure so that alignment focuses on regions near the optimized cuts.
Experimental results show that our approach outperforms state-of-the-art text-to-texture approaches both qualitatively and quantitatively. Qualitatively, our approach is much more photorealistic and does not present discontinuous or blurry texture elements that are frequent among other approaches. Quantitatively, our approach achieves a better FID score than the existing methods.
@@ 2 Related Work
Consistent Multi-View Generation. A challenge in text-to-textured-mesh models is ensuring the multi-view consistency among the text-to-image outputs. Dreamfusion [31] optimizes the geometry and color of a NeRF through SDS loss. Zero123 [25] successfully employs a camera transformation as condition to synthesize novel views from an image. TextMesh [45] leverages quartet canonical views for mesh rendering, yet it tent to yield blurry results. Several methods  [48, 50, 39, 26, 49, 38, 55] are trained on 3D rendered datasets [7] to enhance consistency but tend to lose realistic details in generated images. Moreover, most of them are agnostic to depth condition, which is crucial for alignment between 2D images and 3D meshes.
@@
Text-to-Texture Models. Several pioneers  [29, 30, 4] are trained on a semantic CLIP loss, but still struggle to generate high quality textures with text condition. Trained with a large dataset [35], diffusion models [34] can generate high-quality images with rich details from text prompts. On this basis, additional conditions are added to provide 3D structure-aware through rendering of meshes, such as depth maps and inpainting masks [54]. Some methods [2, 33] generate single view 2D images on these conditions and incrementally merge them into one mesh texture. However, these approaches can easily lead to suboptimal results due to accumulated errors and the lack of global perception. Using a latent UV map [28, 1, 19, 12, 27] as a unified representation has demonstrated its effectiveness of increasing global consistency. Since latent pixels compress a patch in RGB images, the decoded RGB pixels cannot guarantee 3D consistency, which leads to the blurriness or inconsistent details in the textures. The other methods [52, 53] finetune a diffusion model on latent UV space to avoid projection, but also generate artifacts with poor UV expansions. Besides, generating PRB materials is also considered as a focus in some work [8, 51, 3]. In contrast, our approach performs alternating optimization of non-rigid alignment among the selected images, as well as joint optimization of the best cuts to guide the non-rigid alignment.
@@
View Selection. The quality of a multi-view representation is highly dependent on the distribution of source 2D views. View selection is related to the view saliency problem, which has been extensively studied in the existing literature [22, 10, 5, 36, 23, 18, 41]. Closely related to our work, Sun et al. [43] present an approach to optimize camera poses to maximize the coverage of an input scene. Wei et al. [47] optimize viewpoints to maximize the quality of fusion of the learned 2D features. We introduce a view selection approach that selects a subset of consistent views. Driven by the goal of enforcing multi-view consistency, we propose a novel formulation that optimizes a view-consistent score among the chosen images, while constraining that the selected images provide complete coverage of the underlying 3D model.
@@
Texture Cutting and Stitching. Cutting and stitching overlapping images is a fundamental approach to texture synthesis. The pioneering work of Efros and Freeman [11] optimizes cuts between pairs of overlapping patches for texture synthesis. Our method formulates cut optimization and texture stitching as a second-order MRF problem on the mesh faces, where the second-order term models view consistency across different images. This MRF formulation is introduced by Waechter et al. [46], and our contribution lies in the jointly aligned images as input. In particular, we find it essential to alternate between solving the MRF problem and using the MRF output to refine its inputs.
@@ 3.2 Approach Overview
View selection. The second stage selects a subset of RGB-D images that are consistent with each other. This is formulated as a solution to a quadratic assignment problem in which the objective function models the consistency scores between pairs of RGB-D images, under the constraint that the selected images provide a complete coverage of $\mathcal{M}$. We introduce a novel sequential semidefinite programming approach to solve this problem accurately and efficiently.
View alignment. The third stage performs joint non-rigid warping of the selected images so that the pixel colors after warping agree with each other. We first solve a joint optimization problem to alleviate illumination changes in the selected images. We then use SIFTFlow [24] to precompute dense correspondences between overlapping views. The resulting correspondences are fed into another joint optimization problem to optimize the deformation of each view so that the deformed images are view-consistent.
@@ 4 Approach
As illustrated in Fig. 2, our approach consists of four stages. We first describe view generation and view selection in Sec. 4.1 and Sec. 4.2, respectively. We then elaborate on view alignment in Sec. 4.3. Finally, we introduce texture cutting and stitching and discuss how to alternate between this stage and view alignment in Sec. 4.4.
@@ 4.1 Stage I: View Generation
The second phase, which runs between diffusion steps 800-200, adopts latent UV mapping [28], a popular data-free approach to enforce multi-view consistency. The key idea is to aggregate multiple views in a shared latent map that is then mapped back to individual views. Although improved consistency is achieved, image details can still be dropped during the aggregation step, which performs some type of averaging (see supp. material). Therefore, similar to SyncMVD [27], we do not perform latent UV mapping to generate the final output.
@@
As shown in Fig. 3, the multi-view image generation results are consistent in global style and shape. However, local shapes and styles may differ and there are also drifts. Another important idea is to generate an over-complete set of images and operate on a subset of more consistent images to generate the output. This approach nicely mitigates inconsistency introduced in the third phase, which is performed on each view independently.
@@ 4.2 Stage II: View Selection
$\underset{\boldsymbol{z}\in\{0,1\}^{n}}{\textup{minimize}}\quad\boldsymbol{s}^{T}\boldsymbol{z}+\lambda\boldsymbol{z}^{T}S\boldsymbol{z}\qquad\textup{subject to}\quad C\boldsymbol{z}\geq\boldsymbol{1}.$ (1)
Equation 1 is hard to optimize, due to non-convexity and integer variables. We introduce two relaxations. The first relaxes $z_{i}$ so that they can take real values between 00 and $1$, and the second uses the semidefinite programming relaxation to $\boldsymbol{z}^{T}S\boldsymbol{z}$. More precisely, we introduce an additional matrix variable $Y=\boldsymbol{z}\boldsymbol{z}^{T}$ and apply the standard relaxation approach as $Y\succeq\boldsymbol{z}\boldsymbol{z}^{T}$ and $\textup{diag}(Y)=\boldsymbol{z}$. Furthermore, we maintain an index set $\mathcal{I}_{f}$ that specifies all currently selected images. Initially, we $\mathcal{I}_{f}=\emptyset$. This leads to the following semidefinite program:
$\underset{\boldsymbol{z}\in\mathbb{R}^{n},Y\in\mathbb{R}^{n\times n}}{\textup{minimize}}$ $\quad\langle\boldsymbol{s},\boldsymbol{z}\rangle+\lambda\langle S,Y\rangle$
@@
where $Y\succeq\boldsymbol{z}\boldsymbol{z}^{T}$ is identical to Eq. 4. We then augment $\mathcal{I}_{f}$ with $i^{\star}=\textup{argmax}_{i\notin\mathcal{I}_{f}}z_{i}$. This procedure is iterated until $C\boldsymbol{z}\geq 1$ is satisfied. For simplicity, we assume that the selected views are the first $k<n$ images $I_{i},1\leq i\leq k$.
Image score $s(I)$. Our key observation is that if the quality of an image $I$ is high, then the pixel colors tend to repeat among other input images. This motivates us to build a probabilistic color model for each vertex $v\in\mathcal{V}$ (cf., Fig. 4 (Left)). Specifically, denote $\mathcal{I}_{v}\subset\mathcal{I}$ as the subset of images in which $v$ is visible. Let $\boldsymbol{c}(v,I)$ be the color of $v$ in $I$. We first perform mean shift clustering [6] to extract the largest cluster of $\{\boldsymbol{c}(v,I)|I\in\mathcal{I}_{v}\}$, which represents the most represented color. With $\boldsymbol{c}_{v}$ and $\sigma_{v}$ we denote the center and variance of this cluster. We then define
@@
Given an image $I$, let $\mathcal{V}_{I}\subset\mathcal{V}$ collect all visible vertices in $I$. We sort the values $\{p(v,\boldsymbol{c}(v,I))|v\in\mathcal{V}_{I}\}$ in increasing order and set the score $s^{\prime}(I)$ of $I$ as the $t\%$ percentile ($t=20$) of $p(v,\boldsymbol{c}(v,I))$. The final score is given by $s(I)=(s_{\max}-s^{\prime}(I))/(s_{\max}-s_{\min})$ where $s_{\max}(s_{\min})=\max_{I}(\min_{I})s^{\prime}(I)$. As shown in Fig. 4 (Left), high-quality images tend to have small values in $s(I)$.
Consistency score $s(I_{i},I_{j})$. For the purpose of view selection, we find that defining $s(I_{i},I_{j})$ based on mean color differences without alignments is sufficient.
We set $s(I_{i},I_{j})=0$ if $I_{i}$ and $I_{j}$ do not overlap. Figure 4 (Right) shows that the consistency score is critical to selecting consistent images.
@@ 4.3 Stage III: View Alignment
The third stage of our approach performs non-rigid alignment among the selected RGB-D images, making them consistent among their overlapping regions. As illustrated by Fig. 5, we achieve this goal in two steps. The first step adjusts the pixel colors together to reduce the illumination variations in the selected images. The second step performs joint image warping to align image features across the selected images. In both steps, we parameterize the deformation of each image using the free-form deformation (FFD) [37] on a regular $L\times L$ lattice (we choose $L=20$ for color adjustment and $L=40$ for image warping). Note that it is possible to avoid illumination variations by training text-to-image models from rendered 3D shapes. However, our goal is to use photorealistic images from pre-trained text-to-image models. Therefore, we chose to solve an optimization problem to consistently reduce color variations.
@@
$\hat{I}_{i}^{x}(\boldsymbol{p})=B(\boldsymbol{p})\boldsymbol{c}_{i,x}^{1}+I_{i}^{x}(\boldsymbol{p})B(\boldsymbol{p})\big{(}\boldsymbol{1}-\boldsymbol{c}_{i,x}^{0}-\boldsymbol{c}_{i,x}^{1}\big{)},$ (6)
@@
$\min_{\{\boldsymbol{c}_{i,x}^{l}\}}$ $\sum\limits_{x\in\{r,g,b\}}\Big{(}\sum\limits_{i\neq j}\sum\limits_{\boldsymbol{p}\in I_{ij}}\|\hat{I}_{i}^{x}(\boldsymbol{p})-\hat{I}_{j}^{x}(g_{ij}(\boldsymbol{p}))\|^{2}+\mu\sum\limits_{i=1}^{k}\sum\limits_{l=0}^{1}{\boldsymbol{c}_{i,x}^{l}}^{T}L{\boldsymbol{c}_{i,x}^{l}}\Big{)}$
@@
Figure 5: Illustration of the view alignment stage. We show the result of two overlapping images through the joint view alignment procedure. (a) Two input images. (b) After adjustment of the global illumination. (c) From left to right: overlapping regions of each input image, warped images using ground-truth maps between the overlapping regions, overlaid results of the first two columns, and overlaid after SIFTFlow alignment. (d) Overlaid alignments after joint alignment. The pairwise SIFTFlow alignments are preserved.
Image warping. Let $(O_{ij}\subset I_{ij},O_{ji}\subset I_{ji})$ be the focused overlapping subsets between $I_{i}$ and $I_{j}$. Our goal is to deform each $I_{i}$ so that $O_{ij}$ and $O_{ji}$ are aligned among the deformed images. Initially, $O_{ij}=I_{ij}$ and $O_{ji}=I_{ji}$. After determining the optimal cut between $I_{i}$ and $I_{j}$, we recompute $O_{ij}$ and $O_{ji}$ as the dilated regions along the cuts. This allows us to focus on alignments among potentially more consistent regions to improve alignment results.
Let $f_{ij}:g_{ij}(O_{ji})\rightarrow O_{ij}$ and $f_{ji}:g_{ji}(O_{ij})\rightarrow O_{ji}$ be dense pixel-wise alignments between the focused overlapping regions under the view of each image. In our experiments, we compute $f_{ij}$ and $f_{ji}$ using SIFTFlow [24], a state-of-the-art method that establishes high-fidelity dense correspondences (Fig. 5 (c)).
@@
where $\mu_{1}=2$ and $\mu_{2}=200$ in all our experiments. Since Eq. 8 consists of non-linear least squares, we apply the Gauss-Newton method for optimization.